OpenStudy (anonymous):

Can someone explain Simplifying Square Roots with Variables??? I really don't get it.... :/

OpenStudy (anonymous):

Certainly. Do you have an example or two you'd like to look at?

OpenStudy (anonymous):

\[3\sqrt{45x}\] Thanks. :)

OpenStudy (anonymous):

Ok, well the trick here is to factor the part under the radical. Then pull out any squares.

OpenStudy (anonymous):

any question given to help with ?

OpenStudy (anonymous):

So to factor 45x we get?

OpenStudy (anonymous):

kkk miming

OpenStudy (anonymous):

5*9*x right?

OpenStudy (anonymous):

Correct. And \(9 = 3^2\)

OpenStudy (anonymous):

oh, right...

OpenStudy (anonymous):

So we can take the square root of that factor and bring it outside as a multiple

OpenStudy (anonymous):

insert x in the square root in each one of the expression sorry i forgot to put x

OpenStudy (anonymous):

\[3\sqrt{45x}\]\[=3\cdot \sqrt{5\cdot 9 \cdot x}\]\[= 3\cdot \sqrt{5 \cdot 3^2 \cdot x}\]\[=3\cdot 3 \cdot \sqrt{5\cdot x}\]\[=9 \sqrt{5x}\]

OpenStudy (anonymous):

\[3\sqrt{9*5*x}\] \[3*3\sqrt{5x}\] \[9\sqrt{5x}\]

OpenStudy (anonymous):

oh, okay... I get it.

OpenStudy (anonymous):

MIMING 97 have u got it now JUST REMEMBER ::::::::::: \[\sqrt{ab}=\sqrt{a} * \sqrt{b} \]

OpenStudy (anonymous):

YES, THANKS.